OK my word problem is "Melissa is 24 years younger than Joyce. In 2 years, Joyce will be 3 times as old as Melissa will be then. How old are they now?"
My teacher told me to set it up into a table. How i think the table will be set up like this:
Now -------- In 2 yrs
Melissa: x-24 x-24+2
Joyce: x 3(x-24+2)

And when i set up the equation it came out to be:
x-24+2=3(x-24+2)
Im not sure if that is the right way to set it up. I am horrible at Algebra and i really need your help
So my question is: Is my equation set up right or do i need to do it a whole dfferent way?

Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(10682)

You have the right idea!..Just using more info than you need to solve at first.
The way to look at it is that Joyce will "always" be 24 yrs older than Melissa!
so..the 2 years is not important information to set up the equation you need AT FIRST if you just let J be Joyces age in two years time.
THE QUESTION STATES THEN THE FOLLOWING TO BE TRUE(in two years)
Joyce will be 3 times as old as Melissa and, as always, 24 years older.
J =3(J-24)
Simplify:
J = 3*J -72
0=2*J -72
72=2*J
J=36 (THAT'S THE AGE JOYCE WILL BE IN 2 YEARS) Melissa(36/3)or 12 in two years
How old are they now?" Joyce is 34 and Melissa is 10
Check your answer: Joyce is always 24 years older than Melissa.
In this case it was the easiest to just set up what the problem said would be true in two years and work backwards from there!



Answer by Alan3354(30993)

"Melissa is 24 years younger than Joyce. In 2 years, Joyce will be 3 times as old as Melissa will be then. How old are they now?"
My teacher told me to set it up into a table. How i think the table will be set up like this:
Now -------- In 2 yrs
Melissa: x-24 x-24+2
Joyce: x 3(x-24+2)
And when i set up the equation it came out to be:
x-24+2=3(x-24+2)
Im not sure if that is the right way to set it up. I am horrible at Algebra and i really need your help
So my question is: Is my equation set up right or do i need to do it a whole dfferent way?
--------------------
--------------------
M = Melissa's age, J = Joyce's age
J = M + 24 (always true)
J + 2 = 3(M+2) (3x as old in 2 years)
Sub M+24 for J in the 2nd eqn
M+24 + 2 = 3M + 6
-2M + 26 = 6
-2M = -20
M = 10
J = 34
-------
The way you were going might work, but it was giving me a headache.